sam (.sam.):

Math Quiz: A country currency consists of the following coins 1¢, 2¢, 5¢, 10¢, 25¢, 50¢. What is the most money you can have in coins and not be able to pay exactly $1? It can be over a dollar

OpenStudy (anonymous):

Coin change problem.

OpenStudy (unklerhaukus):

i guess you could have 99 1¢ pieces

OpenStudy (anonymous):

This doesn't seem well-posed.

sam (.sam.):

It can be over a dollar

OpenStudy (anonymous):

Then.. There is no limit on what you can pay.

OpenStudy (anonymous):

Or perhaps I am not understanding correctly. You can have \(100\) 1 cent pieces, \(1000\) 1 cent pieces, \(10,000\) 1 cent pieces etc.

OpenStudy (unklerhaukus):

one 50¢ and three 20¢ maybe

OpenStudy (anonymous):

OH. I completely misread the sentence. My bad, yo.

OpenStudy (lgbasallote):

ohhh pop quiz *_*

OpenStudy (anonymous):

you can't hv 2 50's, 4 25's, 1 50's and 2 25's, 1 50's and 5 10's, 1 50 and 20 5's and so on....

OpenStudy (unklerhaukus):

oh there are no 20¢ s only 25¢

OpenStudy (unklerhaukus):

3×25¢ and 3×10¢ =$1.05

OpenStudy (unklerhaukus):

*+4¢

OpenStudy (unklerhaukus):

$1.09

OpenStudy (anonymous):

ans. is 124 cents?

sam (.sam.):

@UnkleRhaukus nope :) , Hint: its greater than that

OpenStudy (anonymous):

50, 25, {10,10,10,10,10,10,10,10,10}, 5, {2,2} ?

OpenStudy (anonymous):

oops, nope.

OpenStudy (anonymous):

First at most 1 50 cents, since 2 50 cents coins = $1 => $0.5 Then, at most 1 25 cents coin , since 2 25 cents coins = $0.5 => + 0.5 above = 1 (=> rejected) Now, we've got $0.75 We can at most 4 $0.1 coins. since 5 x $ 0.1 = $ 0.5 + 0.5 above = $1 (=> rejected) Now, we have $ 0.75 + $ 0.4 = 1.15 We CAN'T have $0.05 coin, since it 0.05 + 0.75 = 0.8 + 0.2 = 1 (=> rejected) we can have at most 4$0.02 coins, since 5x^0.02 = $0.1 and 0.1+ 0.4 = 0.5 (rejected) Now we have 1.23 No 0.01 ... 1.23?

sam (.sam.):

@RolyPoly is correct :D

OpenStudy (anonymous):

Yay!!!~

OpenStudy (anonymous):

50,25, 10, 10,10,10,2,2,2,2? oops 123

OpenStudy (anonymous):

*No 0.01 since 0.01 + 0.02 + 0.02 = 0.05, that is similar to 0.05 case (=> rejected)

OpenStudy (anonymous):

Thanks for the nice problem, .Sam.

sam (.sam.):

np :)

OpenStudy (anonymous):

typo: we can have at most 4$0.02 coins, since 5 x 0.02 = $0.1 and 0.1+ 0.4 = 0.5 (rejected)